A rental car costs $25 plus afixed charge per mile driven. The total chrge for 210 miles of use was $89. Write an equation for the cost,C(in dollars), in terms of the mile driven x.
Answer by uma(370)

The total charge for 210 miles of use was $89.
Let the fixed charge per mile driven be $p
Then the equation is $25 + 210p = 89
==> 210p = 89 - 25 [subtracting 25 from both the sides]
==> 210p = 64
==> 210p/210 = 64/210
==> p = 0.31
so the fixed charge per mile driven = $0.31
The required equation for x miles driven would be
C = 25 + 0.31x